# KnowledgeNavigator: Leveraging Large Language Models for Enhanced Reasoning over Knowledge Graph

### 26 Dec 2023 v1

@@ 4.1 Dataset
To test the ability of KnowledgeNavigator on multi-hop knowledge graph reasoning tasks, we evaluate KnowledgeNavigator on two datasets: MetaQA Zhang et al. (2018) and WebQSP Yih et al. (2016). In the KGQA task on both datasets, we use Hits@1 as the evaluation metric to evaluate the correctness of the answer generated by LLM.
MetaQA is a large-scale KGQA dataset in the movie domain, which provides a knowledge graph with 43k entities, 9 relations and 135k triples, as well as 407k questions. The question set is extracted from the Facebook MovieQA dataset, containing questions that require 1-hop to 3-hop reasoning away from the head entities. Each question contains a head entity, a relation reasoning path, and the answer entities. To verify the multi-hop reasoning ability of KnowledgeNavigator, we choose the 2-hop and 3-hop vanilla datasets in MetaQA.
WebQSP is a benchmark with fewer questions but a large-scale knowledge graph, which can effectively evaluate the large-scale search ability of KnowledgeNavigator. WebQSP provides questions up to 2 hops based on freebase, each question contains topic entity, constraints, inferential chains and SPARQL queries for it’s answer. We set up the knowledge graph with the latest version of freebase data dumps provided by google including 3.12B triples Google (2023). WebQSP provides 4737 questions and we remove 11 questions which have no gold answers from it.
@@ 4.2 Baselines
To evaluate the effectiveness of KnowledgeNavigator in reasoning on knowledge graph, we use a set of well-known baseline models in KGQA field, which are all built with fully supervised for comparison. These baselines include KV-Mem Xu et al. (2019), GraftNet Sun et al. (2018), EmbedKGQA Saxena et al. (2020), NSM He et al. (2021), UniKGQA Jiang et al. (2022). All of these baselines were evaluated on both MetaQA and WebQSP. In addition, we add StructGPT Jiang et al. (2023) and TOG Sun et al. (2023) as baselines of using LLM for KGQA tasks. These two frameworks are both based on un-fine-tuned LLM for knowledge graph retrieval and question reasoning. We apply the results of TOG using the same LLM as KnowledgeNavigator as reference, TOG was only evaluated on WebQSP. Finally, we apply the LLama-2-70B-Chat Touvron et al. (2023b) and ChatGPT as large language model baselines. Specifically, we deploy the same template to prompt the large language models, ensuring that the only difference between this baseline and KnowledgeNavigator is the effectively retrieval knowledge.
@@ 4.3 Experiment Details
KnowledgeNavigator is decoupled from LLM, any LLM can be used as a pluggable component for reasoning. We use ChatGPT and LLama-2-70B-Chat as the LLM component in experiments. LLama-2-70B-Chat is deployed locally with 4 NVIDIA A100 80G with vllm framework Kwon et al. (2023). The context length is set to 4096 as default, and the maximum number of tokens to generate per output sequence is set to 1024. We fine-tune Bert-base-uncased and a linear classifier on the training set of the datasets for hop prediction.
For both datasets, KnowledgeNavigator generates two variants for each question, and the hop prediction is conducted within the range of 1 to 3. In MetaQA, KnowledgeNavigator performs weighted ranking of the top-1 relation for each (question variant, entity, relations) group and selects the top-1 ranked result for the next iteration. For WebQSP, these two parameters are set to top-2. In the few-shot scenario, few-shot is composed of two examples from the training set of the same dataset, which have the same format as the target task.
@@ 4.4 Main Results
First, LLM can answer questions in KGQA tasks without relaying on external knowledge, and even achieves higher performance on WebQSP than KV-Mem. This demonstrates that the knowledge learned by LLM in pre-training can help it answer questions that have not been learned directly. However, the performance of LLM on KGQA tasks still has a large gap with the SOTA models, indicating it is difficult for LLM to perform reasoning and answer complex questions with its own knowledge.
We found that KnowledgeNavigator can support LLM to perform more accurate reasoning and question answering. With KnowledgeNavigator searching and prompting, the performance of LLM on KGQA tasks can be significantly improved, demonstrating its effectiveness. In addition, the performance of KnowledgeNavigator can be further improved by adding a few incontext examples. KnowledgeNavigator achieves the best performance on WebQSP comparing with all baselines, and also outperforms other researches using the same backbone LLM for knowledge retrival and reasoning, such as TOG and StructGPT, on MetaQA. Comparing with models trained with full supervision, KnowledgeNavigator can also achieve comparable performance on MetaQA. KnowledgeNavigator exhibits improvements over KV-Mem with increases of 16.8%, 46.1% and 36.8% respectively on the three datasets. In addition, the performance of KnowledgeNavigator on MetaQA 3-hop is inferior to NSM and UniKGQA. This is due to the complex semantics of multi-hop questions bringing challenges to relation filtering, and also indicates that the deep reasoning ability of LLM still needs to be further improved.
@@ 4.5 Ablation Study
We perform an ablation study on KnowledgeNavigator, aiming to analyze the impact of similar questions and knowledge representation forms. The ablation study includes experiments with different number of similar questions and knowledge representation forms. In the ablation study, we use LLama-2-70B-Chat as the backbone LLM and use the same prompt template with 2-shot examples for all cases. Table 2 and 3 shows the results of the ablation study.
@@ 4.5.1 Impact of Number of Similar Questions
Table 2 shows the accuracy of using 0, 2, and 4 similar questions for voting in relation selection. For MetaQA, the performance of KnowledgeNavigator increases slowly with the number of similar questions, since more voters make the relation selection more stable. For WebQSP, the low quality of the original questions makes it difficult for LLM to generate similar questions with exactly the same semantics in some cases, so the similar questions have limited help to KnowledgeNavigator and may cause voting errors. In addition, since the number of requests of KnowledgeNavigator to LLM increases linearly with the number of similar questions, there is a trade-off between the computational cost and the performance of KGQA. For the experiments, we use 2 similar questions as default to control the computational cost.
@@ 4.5.2 Impact of Knowledge Formats
Table 3 shows the impact of different knowledge representation forms on the performance of KnowledgeNavigator. In this part, we use the same knowledge with different representation forms to prompt LLM. Specifically, for “w/ single triples” and “w/ joined triples”, all triples are concatenated into a string in the form of $[head,relation,tail]$ or $[head,relation,[tail_{1},...,tail_{n}]]$, for “w/ single sentences”, each triple is converted into a separate natural language sentence using a template and concatenated into a string.
For different knowledge formats, we found that the performance of KnowledgeNavigator increases with the logical closeness of the knowledge representation. First, for both triples and sentences, using aggregated knowledge can effectively reduce the redundant information and improve the density of the knowledge in the prompt, therefore reduce the difficulty of LLM reasoning. Second, for the general LLM without fine-tuning, using knowledge in natural language format can avoid the errors caused by the insufficient understanding ability of LLM on the structure or relation information of triple knowledge.
@@ 4.6 Error Analysis
    Other Errors: Including entities answered by large models are incomplete, knowledge truncation due to excessively long context, etc.
Table 4 shows the error analysis results on the three datasets. It is easy to find that the error distribution on the the datasets are different, the reasoning error is the main error on MetaQA 2-hop while the relation selection error appears most on MetaQA 3-hop and WebQSP. This is because the semantics of the questions in MetaQA 3-hop and WebQSP are more complex: Question in MetaQA 3-hop features a longer reasoning path and a more complex structure of knowledge sub-graph, challenging the reasoning capabilities of LLM. Each entity in WebQSP is associated with dozens of relations, many of which are similar to each other, making it extremely difficult for LLM to select the high relevant relations for next iteration. As the reasoning logic of questions in MetaQA 2-hop is more obvious, LLM rarely selects the wrong relations, but shows insufficient reasoning ability.
According to the error statistics, we can improve the performance of LLM on KGQA tasks through targeted optimization. Specifically, we can improve its ability of selecting the best relations by enhancing the semantics of the questions or strengthening the connecting between the knowledge graph and the reasoning path. We can also help LLM to perform reasoning by optimizing the prompt and knowledge representation. We will keep working on these issues in the future.
@@ 4.7 Case Study
Figure 2 is an example of KnowledgeNavigator performing a KGQA tasks. First of all, KnowledgeNavigator predicts the reasoning hops starting from the core entity “Babaloo Mandel” as 2 based on the PLM, and generates 2 similar questions for the target question with LLM.
Then, KnowledgeNavigator extracts all relations linked to “Babaloo Mandel” and serializes them into “birth_year; birth_place; wrritten_by; created_by”, as part of the prompt. For the only core entity “Babaloo Mandel”, LLM selects the optimal relation linked to it based on each similar question and gets the weighted voting result {written_by:3, created_by:1}. As the number of optimal relation selections for each entity is set to one on MetaQA, the triples with “written_by” as the optimal relation and “Babaloo Mandel” as head or tail entity (i.e. [Splash, written_by, Babaloo Mandel] and [Parenthood, written_by, Babaloo Mandel]) are extracted as the first step of the reasoning path. The tail entities “Splash” and “Parenthood” are selected as the starting entities for the second iteration to continue the knowledge retrieval.
After KnowledgeNavigator iterates to the predicted hops, the triples retrieved from the knowledge graph can be combined into an effective reasoning path (e.g. Babaloo Mandel - written_by - Splash - starred_actors - Dianne Wiest), and further into a reasoning sub-graph. For the knowledge about “Splash”, KnowledgeNavigator will first combine the triples into “[Splash, startted_actors, [Dary Hannah, Tom Hanks]]”, and then convert it into “The actors starred in Splash are: Dary Hannah and Tom Hanks” using a template. All triples retrieved are finally concatenated into a string and feed to LLM as part of the answer generation prompt.
@@ 5 Conclusion
In this paper, we study the knowledge limitation problem in LLM and propose a framework KnowledgeNavigator to improve the reasoning and question answering ability of LLM on knowledge graph. KnowledgeNavigator consists of three stages: Question Analysis, Knowledge Retrieval and Reasoning. KnowledgeNavigator first pre-analyzes the given question and generates similar questions to assist reasoning. Then, relying on the guidance of LLM and the question, it iteratively retrieves and filters candidate entities and relations in the knowledge graph to extracts relevant external knowledge. Finally, it converts the knowledge into an effective prompt to enhance the performance of LLM on knowledge-intensive tasks. We use KGQA metrics to evaluate KnowledgeNavigator, and the experimental results show that introducing external knowledge from the knowledge graph is helpful for LLM on knowledge-intensive tasks. KnowledgeNavigator outperforms other frameworks using knowledge graph to enhance LLM and achieves comparable performance with previous fully supervised models. We also verify the effectiveness of each component in KnowledgeNavigator through ablation study and analyze the errors to show the further research directions of KnowledgeNavigator.
@@ 6 Limitations
Although KnowledgeNavigator has shown impressive performance on KGQA tasks, our method still has some limitations. First, the logic path extraction and knowledge reasoning are based on feeding a prompt built with dozens of sentences to LLM. Therefore, the performance of KnowledgeNavigator is closely related to the natural language understanding and reasoning ability of the backbone LLM. It may be difficult to generalize KnowledgeNavigator to LLM with less parameters or shorter context length. Second, the answer generation process of KnowledgeNavigator contains multiple interactions with LLM, which leads to a long delay especially in questions with complex reasoning paths and may not be suitable for scenarios with high response time requirements. Therefore, further research should focus on these limitations to improve the application of LLM in KGQA.

### 19 Jan 2024 v2

@@ 4.1 Dataset
To test the ability of KnowledgeNavigator on multi-hop knowledge graph reasoning tasks, we evaluate it on two datasets: MetaQA Zhang et al. (2018) and WebQSP Yih et al. (2016). In the KGQA task on both datasets, we use Hits@1 as the evaluation metric to evaluate the correctness of the answers generated by LLM.
MetaQA is a large-scale KGQA dataset in the movie domain, which provides a knowledge graph with 43k entities, 9 relations and 135k triples, as well as 407k questions. The question set is extracted from the Facebook MovieQA dataset, containing questions that require 1-hop to 3-hop reasoning away from the head entities. Each question consists of a head entity, a relation reasoning path, and the answer entities. To verify KnowledgeNavigator’s multi-hop reasoning capability, we specifically use the 2-hop and 3-hop vanilla datasets in MetaQA for experiment.
WebQSP is a benchmark with fewer questions but a large-scale knowledge graph, which can effectively evaluates the large-scale search ability of KnowledgeNavigator. WebQSP provides questions up to 2 hops based on freebase, each question contains a topic entity, constraints, inferential chains and SPARQL queries for finding the answer. We set up the knowledge graph with the latest version of freebase data dumps provided by google including 3.12B triples Google (2023). WebQSP provides 4737 questions and we remove 11 questions which have no gold answers from it.
@@ 4.2 Baselines
To evaluate the effectiveness of KnowledgeNavigator in reasoning on knowledge graphs, we compare it with a set of well-known baseline models in the field of KGQA, which are all built with fully supervised. These baselines include KV-Mem Xu et al. (2019), GraftNet Sun et al. (2018), EmbedKGQA Saxena et al. (2020), NSM He et al. (2021), UniKGQA Jiang et al. (2022). All of these baselines were evaluated on both MetaQA and WebQSP. In addition, we add StructGPT Jiang et al. (2023) and TOG Sun et al. (2023) as baselines of using LLM for KGQA tasks. These two frameworks are both based on un-fine-tuned LLM for knowledge retrieval and question reasoning. We use the results of these two frameworks using the same LLM as KnowledgeNavigator as a reference. It is important to note that TOG was only tested on WebQSP.
We apply the LLama-2-70B-Chat Touvron et al. (2023b) and ChatGPT as large language model baselines. Specifically, we deploy the same template to prompt the large language models. The only distinction between this baseline and KnowledgeNavigator is the external knowledge retrieved.
@@ 4.3 Experiment Details
KnowledgeNavigator is decoupled from LLM, any LLM can be used as a plug-in component for reasoning. We use ChatGPT and LLama-2-70B-Chat as the LLM component in experiments. We call ChatGPT with the OpenAI API. LLama-2-70B-Chat is deployed locally with 4 NVIDIA A100 80G with vllm framework Kwon et al. (2023). The context length is set to 4096 as default, and the maximum number of tokens to generate per output sequence is set to 1024. We fine-tune bert-base-uncased and a linear classifier on the training set of the datasets for hop prediction.
For both datasets, KnowledgeNavigator generates two variants for each question, and the hop prediction is conducted within the range of 1 to 3. In MetaQA, KnowledgeNavigator performs weighted ranking on the top-1 relation for each (question variant, entity, relations) group and selects the top-1 ranked result for the next iteration. For WebQSP, these two parameters are set to top-2. In the few-shot scenario, few-shot is composed of two examples from the training set of the same dataset, which is in the same format as the target task.
@@ 4.4 Main Results
First, LLM can answer questions in KGQA tasks without relaying on the external knowledge, and even outperforms KV-Mem on the WebQSP benchmark. This demonstrates that the LLM’s pre-training knowledge is useful for answering questions that have not been learned directly. Nonetheless, there is still a significant performance gap between the LLM and state-of-the-art models in KGQA tasks. This suggests that the LLM faces challenges in reasoning and answering complex questions using only its internal knowledge.
We found that KnowledgeNavigator can enhance the accuracy of LLMs in reasoning and question answering tasks. By employing KnowledgeNavigator’s search and prompting capabilities, LLMs achieve significant improvements on KGQA. In addition, the performance of KnowledgeNavigator can be further improved by adding a few incontext examples. In a comparison with all baseline models, KnowledgeNavigator achieves best results on WebQSP. KnowledgeNavigator surpasses other research using the same backbone LLM for knowledge retrieval and reasoning, such as TOG and StructGPT, on MetaQA. Even against fully supervised models, KnowledgeNavigator can also achieve competitive performance on MetaQA. KnowledgeNavigator exhibits improvements over KV-Mem with increases of 16.8%, 46.1% and 36.8% respectively on the three datasets. However, KnowledgeNavigator’s performance on MetaQA 3-hop questions is inferior to NSM and UniKGQA. This is due to the complex semantics of multi-hop questions which bring challenges to relation filtering, and suggest that the deep reasoning capability of LLM still needs to be further enhanced.
@@ 4.5 Ablation Study
We perform an ablation study on KnowledgeNavigator, aiming to analyze the impact of similar questions and knowledge representation forms. The ablation study involves experiments with varying number of similar questions and forms of knowledge representation. In the ablation study, we use LLama-2-70B-Chat as the backbone LLM and use the same prompt template with 2-shot examples for all cases. Table 2 and 3 shows the results of the ablation study.
@@ 4.5.1 Impact of Number of Similar Questions
Table 2 shows the accuracy of using 0, 2, and 4 similar questions for voting in relation selection. For MetaQA, the performance of KnowledgeNavigator increases slowly with the number of similar questions, since more voters make the relation selection more stable. For WebQSP, the low quality of the original questions makes it difficult for LLM to generate similar questions with identical meanings. Consequently, these similar questions have limited help to KnowledgeNavigator and may lead to voting errors. Moreover, as the number of KnowledgeNavigator’s requests to the LLM rise linearly with the number of similar questions, there’s a balance to be struck between computational costs and it’s effectiveness. In our experiments, we use 2 similar questions as default to control the computational cost.
@@ 4.5.2 Impact of Knowledge Formats
Table 3 shows the impact of different knowledge representation forms on the performance of KnowledgeNavigator. In this part, we use the same knowledge with different representation forms to prompt LLM. Specifically, for “w/ single triples” and “w/ joined triples”, all triples are concatenated into a string in the form of $[head,relation,tail]$ or $[head,relation,[tail_{1},\ldots,tail_{n}]]$, for “w/ single sentences”, each triple is converted into a separate natural language sentence using a template and concatenated into a string.
For different knowledge formats, we found that the performance of KnowledgeNavigator increases with the logical closeness of the knowledge representation. First, for both triples and sentences, using aggregated knowledge can effectively reduce the redundant information and improve the density of the knowledge in prompt, therefore reduce the difficulty of reasoning. Second, for the general LLMs that haven’t been fine-tuned, using knowledge in natural language format can avoid the errors caused by their insufficient understanding capabilities on structured data.
@@ 4.6 Error Analysis
    Other Errors: Including intermediate errors causing search interruption, and excessively long context leading to knowledge truncation, etc.
Table 4 shows the error analysis results on the three datasets. It is easy to find that the error distribution on the three datasets are different. Furthermore, the reasoning error is the main error on MetaQA 2-hop and the relation selection error appears most on MetaQA 3-hop and WebQSP. This is because the semantics of the questions in MetaQA 3-hop and WebQSP are more complex: Question in MetaQA 3-hop features a longer reasoning path and a more intricate knowledge sub-graph, challenging the relation reasoning capabilities of LLM. For WebQSP, each entity is associated with numerous similar relations in freebase, which complicates the task for LLM to select the high relevant relations for the next iteration. Meanwhile, as the reasoning logic of questions in MetaQA 2-hop is more straightforward, LLM rarely selects the wrong relations, but shows insufficient reasoning ability.
According to the error statistics, we can improve the performance of LLM on KGQA tasks through targeted optimization. Specifically, we can improve its ability of selecting the relevant relations by enhancing the semantics of the questions or strengthening the connecting between the knowledge graph and the reasoning path. We can also help LLM to perform reasoning by optimizing the prompt and knowledge representation. We will keep working on these issues in the future.
@@ 4.7 Case Study
Figure 2 is an example of KnowledgeNavigator performing a KGQA task. First of all, KnowledgeNavigator predicts the reasoning hop number starting from the core entity “Babaloo Mandel” based on a PLM, and generates 2 similar questions for the target question with LLM.
Then, KnowledgeNavigator extracts all relations linked to “Babaloo Mandel” and serializes them into “birth_year; birth_place; wrritten_by; created_by”, as part of the prompt. For the only core entity “Babaloo Mandel”, LLM selects the optimal relation linked to it based on each question variant and gets the weighted voting result {written_by:3, created_by:1}. As the number of optimal relation selection for each entity is set to one, the triples with “written_by” as the optimal relation and “Babaloo Mandel” as head or tail entity (i.e. [Splash, written_by, Babaloo Mandel] and [Parenthood, written_by, Babaloo Mandel]) are extracted as the first step of the reasoning path. The tail entities “Splash” and “Parenthood” are selected as the core entities for the second iteration to continue the knowledge retrieval.
After KnowledgeNavigator reached the predicted hops, the triples retrieved from the knowledge graph can be combined into an effective reasoning path (e.g. Babaloo Mandel - written_by - Splash - starred_actors - Dianne Wiest), and further into a reasoning sub-graph. Take the knowledge about “Splash” as example, KnowledgeNavigator will first combine the triples into “[Splash, startted_actors, [Dary Hannah, Tom Hanks]]”, and then convert it into “The actors starred in Splash are: Dary Hannah and Tom Hanks” with template. All triples retrieved are finally concatenated into a string and feed to LLM as part of the answer generation prompt.
@@ 5 Conclusion
In this paper, we study the challenge of knowledge limitations in LLM and introduce a framework KnowledgeNavigator to improve the reasoning and question answering capabilities of LLM on knowledge graph. KnowledgeNavigator consists of three stages: Question Analysis, Knowledge Retrieval and Reasoning. During Question Analysis, KnowledgeNavigator first pre-analyzes the question and generates variants for it to assist reasoning. Then, relying on the guidance of LLM, it iteratively retrieves and filters candidate entities and relations within the knowledge graph to extracts relevant external knowledge. Finally, this knowledge is transformed into an effective prompt to improve LLM’s performance on knowledge-intensive tasks. We evaluate KnowledgeNavigator with KGQA metrics, and the results indicate that introducing external knowledge from the knowledge graph benefits LLM in handling complex tasks. KnowledgeNavigator outperforms other frameworks deploying LLM on enhanced KGQA and achieves comparable performance with previous fully supervised models. We also conduct an ablation study to confirm the effectiveness of each component in KnowledgeNavigator and analyze the errors to identify the further research directions.
@@ 6 Limitations
Although KnowledgeNavigator has shown impressive performance on KGQA tasks, it still has some limitations. First, the logic path extraction and knowledge reasoning are achieved by feeding prompts built with multiple sentences to LLM. Therefore, the performance of KnowledgeNavigator is closely related to the natural language understanding and reasoning ability of the backbone LLM. It might be challenging to apply KnowledgeNavigator to LLMs with fewer parameters or shorter context lengths. Second, the answer generation process of KnowledgeNavigator contains multiple interactions with LLM. This can result in longer delays, especially in questions with complex reasoning paths, and may not be suitable for scenarios that have strict response time requirements. Therefore, further research should focus on these limitations to improve the application of LLM in KGQA.